OpenStudy (anonymous):

It takes Karina 3.5 hours to drive from Davenport to Hogsdale and then 2 hours and 10 minutes to drive from Hogsdale to Paqua. On the route that she took, it is 720 miles from Davenport to Paqua. What was her average driving speed for the trip?

OpenStudy (calculusxy):

3.5x=720 3.5 is 350/100 The reciprocal of 350/100 is 100/350. So 720/1*100/350=72000/350=2057.1428

OpenStudy (anonymous):

that confused me more lol. but what happen to the other 2hrs & 10 mins, wouldnt the total hours be 5hrs & 40 minutes?